# From attention to profit: quantitative trading strategy based on transformer

### v1

@@ 3.2 Problem formulation
For a stock $s_{k}^{t}$ in stock set $S^{t}$, where $k\in N$, it details a sequence for a historical 1-time step (such as one month), and the dimension is 2:
@@
In each $x_{k}^{t}$, there are two features. The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ is the close price of the $i^{th}$ and ${i-1}^{th}$ day, respectively.
@@
Another is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
@@
For a stock set $S^{t}$ on the trading timestamp $t$, each stock $S^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is same as the method of $r^{t}$ in the sequence of inputs shown in Eq.(1). Then the value in the list of next-time stamp’s profit for $N$ stocks in the timestamps $t+1$ can be represented as:
@@
And the list will be arranged as $n$ parts and each part has an equal proportion $\frac{1}{n}$ ($n$ should be an odd number, such as 3,5,7). Take $n=5$ as an example. For those ranked values in the first, middle and last $\frac{1}{5}=20\%$, the output of the stock $s_{k}$ records as below, respectively:
$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
@@ 3.3 Embedding initialization
The next step is the normalize the input data for each time step by zero-mean normalization (Z-Score), the normalization equation is shown in Eq.(2).
$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{th}$ day of stock $k$ in time $t$.
$E[x^{t}_{i}]$ and $std[x^{t}_{i}]$ is the mean and standard of all the value of $x^{t}_{i}$, respectively.
$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-E[x^{t}_{i}]}{std[x^{t}_{i}]}$ (2)
In this way, the input consequences towards zero mean and unit variance, which aims to reduce the influence from different time points and allow different features to be comparable with each other [39].
3.4 Transformer encoder
Then a transformer encoder structure will be built in this subsection. The representation hierarchy consists of $L$ blocks of multi-head self-attention layers. Taken initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, and the canonical self-attention in [27] performs information exchange between every time points for each stock $s_{i}$.
To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding. On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value or the probability of price increase or decrease. Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations.
For each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$, transforms into query, key, and value metrics, respectively:
$\textbf{Q}_{i,h}=\textbf{E}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{E}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{E}_{i}\textbf{W}^{V}_{h}$ (3)
where $h=1,\ldots,H$ is the head index, as well as $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$. Lastly, the final layer is represented by the concatenation of all attention heads.
$\textbf{F}_{i}=Multihead(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=||^{h=H}_{h=1}Softmax(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h})$ (4)
Just like Eq.(4) shows, ||||| | represents the concatenation operator, and $d_{f}$ is the dimension of projected feature space. In the end, all input consequences are represented in the form of:
@@ 3.4 Transformer encoder
which is followed by feed-forward layers.
According to Sec.3.2, the embedding output is denoted as:
$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$
@@ 3.5 Output process
Eventually, the output is represented:
$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$ (5)
The output will be activated by the $softmax$ function. And for each $z^{t}_{i}$, there are:
$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,n}]\in\mathbb{R}^{1\times n}$
the output of probability can be represented as (m is the middle number from 1 to n):
$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,n}]^{T}=[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{i}}}]$ (6)
In Eq.(6), $y^{t}_{o,i}$ are larger than 0 and smaller than 1 as well as the sum of all the items in $Y^{t}_{o}$ is 1. So the output means the probability of the stock’s raising, flatting, or downing in the next timestamp.
@@ 3.6 Prediction
The Mean Squared Error Loss (MSELoss) will be used to quantify the loss. MSE Loss is a widely used metric for assessing the discrepancy between a model’s predictions and the actual values. It calculates the average of the squares of the differences between the predicted value and the actual prices. This is formally represented as:
$MSELoss=\frac{1}{n}\sum^{n}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (7)
In Eq.(7), $y_{i}$ is the actual value of the $i^{th}$ equation, $\hat{y}_{i}$ is the model’s prediction for the same, and n represents the total number of samples in the inputs. During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this value over successive iterations. A lower MSE indicates a closer alignment between the model’s predictions and the actual stock prices, signifying an improvement in the model’s learning and its ability to generalize from the training data. This loss function is particularly useful in regression tasks, as it not only penalizes the model for inaccurate predictions but also emphasizes reducing larger errors, thereby ensuring a robust learning process.
@@ 4.1 Dataset
The training data of the transformer factor comes from the Chinese exchange market. The dataset is collected from the Shanghai Stock Exchange (SHSE) and Shenzhen Stock Exchange (SZSE), which contains 4601 stocks that are listed or had been listed from January 2010 to May 2023. And all these data come from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which is one of the quantitative finance terminals. The train period is from January 2010 to December 2019 and the test period starts from January 2020.
Adjustment closing price is used for the stock training price [40]. Adjustment closing price is essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns. This price takes into account factors like dividends, stock splits, and other corporate actions that can affect a stock’s price over time [41]. By adjusting for these events, the adjusted close price provides a more accurate picture of a stock’s value and performance, which is more persuadable in quantitative financial back-test.
@@ 4.2 Timestamp
To test the performance of the factor better, three kinds of frequency are chosen, which are stock data monthly, weekly, and daily, respectively.
In the first experiments, a timestamp was set as one month. So for a stock $s_{i}$ from stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month. Both the accumulated profit and accumulated turnover rate are the sum of the trading day’s data in that month. In some cases, if one stock was first listed on the market or resumed trading in one month, the available data may not cover whole trading days, the data of the stock from that from would still be recorded. However, if there is a whole month that the stock did not have the trading record (perhaps stopped trading or not listed in that month), this month’s data will be recorded as ”NaN”. In the sequence $\chi^{t}_{k}$, the sequence will not be used to train the model when there are missing values in the sequence.
Within the first experiments, three sub-experiments were set. For the sub-experiments, the input of them are same. For output data, outputs with different dimensions and different lengths but the same dimension are used. $n=5$ and $n=7$ are used in this experiments. In the group of $n=5$, the first one deleted the data with output like $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
Most of the accumulated profit values in different years are in the range $(-10\%,10\%)$. There are some values point outside the range $(-100\%,100\%)$ that are shown in it because these values are too large (small), such as the stock ”002961.SZ” first listed at SZSE on September $4^{th}$, 2019 and its accumulated profit rate on that month is 184.66%. These stocks received unusual performance mainly because they were first listed on the market, resumed trading, or reported great positive/negative news.
Most of the accumulated turnover rate values are in the range $(0,100\%)$. There are some values outside the range $(0,200\%)$ that are shown in it because these values are too large, such as the accumulated turnover rate in July 2015 of ”002596.SZ” is $777.67\%$, this was because the actual controller abundantly increased his holdings of the company and the company announced that they received findings by shareholders. Similar to the example, the situation may be caused by great positive or negative information.
@@
In the second and third experiments, the inputs are in weekly and daily frequency, respectively. The sub-experiments under different groups are similar to the first one. All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in the Table 1.
@@ 4.3 Implementation Details
The training process and model are implemented with PyTorch [42]. The parameter of the model is searched by gird search, which is reliable in low dimensional spaces and simple to implement [43]. As a result, in the situation of input dimension is 2, the dimension of hidden feature space $d_{f}$ is 16. The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6. The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU by Adam optimizer [44] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
@@ 4.4 Trading strategy
A trading strategy is used for the model and Algorithm (1) shows the algorithm of strategy. Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from stock set $S^{t}$ will be put in the model and obtain the list of the output. Then the stocks will be ranked according to the first element of the output and the first $\frac{1}{n}\%$ stocks will be added to the stock pool. If the stock has already been in the stock pool on the last timestamp, it will keep holding; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account. Stocks that are not in the predicted pool will be sold out. The same method will run repeated during the rest period.
The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade.
@@ 4.5 Metrics
Formally, the Sharpe Ratio (SR) and the $\alpha$ ratio will be used to test the performance of the strategy. The SR [3] is a measure of risk-adjusted return which can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8), where $R_{p}$ is the return of the portfolio, $R_{f}$ is the risk-free rate:
$SR=\frac{E[R_{p}]-R_{f}}{std[R_{p}]}$ (8)
Alpha represents the excess return of an investment portfolio relative to its benchmark. It reflects the stock selection skills of an investment. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment. Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return. Eq.(9) shows the calculation method of alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
@@
5 Result and discussion
Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes. The discussion focuses on the comparative effectiveness of monthly, weekly, and daily strategies. The results underscore the nuanced relationship between trading frequency and market performance, offering valuable insights for the development of robust quantitative trading strategies.

### v2

@@ 3.2 Problem formulation
For a stock $s_{k}^{t}$ in the stock set $S^{t}$, where $k\in\{1,\ldots,N\}$, consider the two-dimensional historical 1-time step sequence (such as one month):
@@
Each $x_{k}^{t}$ contains two features. The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ are the close price of the $i^{\mathrm{th}}$ and ${(i-1)}^{\mathrm{th}}$ day, respectively.
@@
The other feature is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
@@
For a stock set $S^{t}$ on the trading timestamp $t$, each stock $s^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is the same as for $r^{t}$ in the sequence of inputs shown in equation (1). Then the value in the list of next-time stamp’s profit for $N$ stocks in the timestamps $t+1$ can be represented as:
@@
This list is then sorted and partitioned into $q$ equal parts, where it is set to 3 or more. For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively. And the dimension of the outputs $d_{f}$ is the number of selected quantiles in training. The value of $d_{f}$ should be larger than 3 and not larger than $q$.
Figure 4: Training outputs remarked between different $d_{f}$
Both the dimensions of training and predicted outputs are the same, which is the $d_{f}$ set at the beginning. For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively, and the corresponding output of $d_{f}=3$ is recorded, respectively, as
$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
The dimension of the output also can be 5, and in this case, the outputs of $d_{f}=5$ are:
$y^{t}_{k}=[1,0,0,0,0]^{T},\quad y^{t}_{k}=[0,0,1,0,0]^{T},\quad y^{t}_{k}=[0,0,0,0,1]^{T}$
In this example, when $d_{f}=3$, the output dimension will be 3 while the dimension will be 5 as $d_{f}$ is set as 5.
@@ 3.3 Embedding initialization
The next step is to normalize the input data for each time step by zero-mean, unit-variance normalization (Z-score), the normalization equation is shown in equation (2).
$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{\mathrm{th}}$ day of stock $k$ at time $t$. $\mathbb{E}[x^{t}_{i}]$ and $\mathrm{std}[x^{t}_{i}]$ are the mean and standard deviation of the variables $x^{t}_{i}$, respectively.
$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-\mathbb{E}[x^{t}_{i}]}{\mathrm{std}[x^{t}_{i}]}$ (2)
In this way, the input sequences are normalized with zero mean and unit variance, which aims to reduce the influence from outlying time points and allow different features to be comparable with each other [47].
3.4 Quantformer encoder
Then a quantformer encoder structure will be built in this subsection. The representation hierarchy consists of $L$ blocks of multi-head self-attention layers. Taking initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, the canonical self-attention in [10] can perform information exchange between every time points for each stock $s_{i}$.
To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding. On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value representing the probability of price increase or decrease. Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations.
More specifically, each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$ is transformed into query, key, and value metrics, respectively:
$\textbf{Q}_{i,h}=\textbf{P}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{P}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{P}_{i}\textbf{W}^{V}_{h}$ (3)
where $h=1,\ldots,H$ is the head index, as well as trainable weights $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$. Then, the final layer is represented by the concatenation of all attention heads.
$\textbf{F}_{i}=\mathrm{Multihead}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\big{(}||^{h=H}_{h=1}\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})\big{)}\textbf{W}^{O}$ (4)
where ||||| | represents the concatenation operator and $d_{f}$ is the dimension of the projected feature space. After the attention computation is completed, the multi-head attention output is divided into different ”heads”. These outputs from the heads need to be recombined back into the original input dimension. $\textbf{W}^{O}$ $\in\mathbb{R}^{Hd_{f}\times 2}$ in equation 5 plays the role of combining these separate heads back into the original dimension. The attention head is shown below:
$\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\mathrm{softmax}(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h}$ (5)
In the end, all input sequences are represented in the form of:
@@ 3.4 Quantformer encoder
which are fed into feed-forward layers. From Section 3.2, the embedding output is denoted as:
$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$
@@ 3.5 Output process
Eventually, the output is represented as:
$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$ (6)
The output layer uses the softmax activation function, and for each $z^{t}_{i}$ can be written as:
$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,d_{f}}]\in\mathbb{R}^{1\times d_{f}}$
($m=1,2,\ldots,n$ is the index number between $1$ and $n$). The output probability can be represented as:
$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,d_{f}}]^{T}=\left[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{d_{f}}}}\right]$ (7)
In equation (7), the $y^{t}_{o,i}$ take values between 00 and $1$ and sum to $1$, which means that they can be interpreted as the probability of the stock’s performance in the slice $m$ in the next timestamp
@@ 3.6 Prediction
The Mean Squared Error Loss (MSELoss) will be used to quantify the loss. MSE Loss is a widely used metric for assessing the discrepancy between a model’s predictions and the actual values. It is defined as the average of the squares of the differences between the predicted value and the actual prices:
$\mathrm{MSELoss}=\frac{1}{N}\sum^{N}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (8)
In equation (8), $y_{i}$ is the actual value of the $i^{\mathrm{th}}$ observation, $\hat{y}_{i}$ is the corresponding model prediction, and $N$ is the total number of input samples. During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this loss over successive iterations. A lower MSE indicates a closer alignment between the model’s predictions and the actual stock prices, signifying an improvement in the model’s learning and its ability to generalize from the training data.
@@ 4.1 Dataset
The training data of the quantformer comes from the Chinese exchange market. The dataset is collected from the Shanghai Stock Exchange (SHSE) and Shenzhen Stock Exchange (SZSE), which contains 4601 stocks that are listed or had been listed from January 2010 to May 2023. The data comes from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which are quantitative finance terminals. The training period is from January 2010 to December 2019 and the testing period starts from January 2020.
Closing price adjustments are used for the stock training price [48]. Closing price adjustments are essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns. This price takes into account factors such as dividends, stock splits, and other corporate actions that can affect a stock’s price over time [49]. By adjusting for these events, the adjusted closing price provides a more accurate picture of a stock’s value and performance, which is more appropriate for financial backtesting.
@@ 4.2 Timestamp
Data frequency concerns the number of data points within a specific unit of time, which may reflect different market characteristics [50]. To better test the performance of the factor, three frequencies are considered: monthly, weekly, and daily stock data.
In the first experiments, a timestamp was set as one month. So for a stock $s_{i}$ from the stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month. Both the accumulated profit and accumulated turnover rate are the sum of the trading day’s data in that month. In some cases, if one stock was first listed on the market or resumed trading in one month, the available data may not cover whole trading days, the data of the stock from that from would still be recorded. However, if there is a whole month during which the stock did not have any trading (perhaps due to stopped trading or not being listed in that month), this month’s data will be recorded as “NaN”. In the sequence $\chi^{t}_{k}$, the sequence will not be used to train the model when there are missing values in the sequence.
Within the first experiments, three sub-experiments were set. The sub-experiments share the same inputs. For output data, outputs with different dimensions and different lengths but the same dimension are used. $n=5$ and $n=7$ are used in this experiments. In the group of $n=5$, the first one deleted the data with output such as $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
@@
In the second and third experiments, the inputs are in weekly and daily frequency, respectively. The sub-experiments under different groups are similar to the first one. All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in Table 1.
@@ 4.3 Implementation details
The training process and model are implemented with PyTorch [51]. The parameters of the model are optimized by grid search, which is reliable in low dimensional spaces and simple to implement [52]. Here the input dimension is 2, and the dimension of the hidden feature space $d_{f}$ is 16. The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6. The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU using the Adam optimizer [53] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
@@ 4.4 Trading strategy
A trading strategy is used for the model and Algorithm 1 shows the pseudocode of the strategy. Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from the stock set $S^{t}$ will be put in the model and obtain the list of outputs. Then, the stocks will be ranked according to the first element of the output and the first $\frac{1}{q}\%$ stocks will be added to the stock pool. If the stock already was in the stock pool on the last timestamp, it will be held; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account. Stocks that are not in the predicted pool will be sold out. The same method is run repeatedly during the subsequent periods. The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade.
@@ 4.5 Metrics
Formally, the Sharpe Ratio (SR) and the $\alpha$ ratio will be used to test the performance of the strategy. The SR [3] is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9), where $R_{p}$ is the return of the portfolio and $R_{f}$ is the risk-free rate is the rate of London Interbank Offered Rate (LIBOR) which calculated averaged from estimates submitted by banks in London. The risk-free rate in the backtest will be calculated under the average LIBOR rates during the test period.
$\mathrm{SR}=\frac{\mathbb{E}[R_{p}]-R_{f}}{\mathrm{std}[R_{p}]}$ (9)
Alpha represents the excess return of an investment portfolio relative to its benchmark. It reflects the stock selection skills of an investment. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment. Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return. Equation (10) shows the calculation method of the alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
@@
$VaR=\mu+\sigma N^{-1}(X)$ (11)
[…]
5 Results and discussion
Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes. The discussion focuses on the comparative effectiveness of different timestamps (monthly, weekly, and daily) and different training scales. The results underscore the nuanced relationship between trading frequency and market performance, offering valuable insights for the development of robust quantitative trading strategies.